OpenStudy (anonymous):

Is my answer right??? Directions: State wether the percent of change is a percent of increase or decrease, Then find the percent of change. Round to the nearest whole percent. Formula: Final-original/original=r/100 Original-11 New-33 22/11=r/100 22(100)=11(r) 2200=11r /11 /11 r=200 Increase, 200%

hero (hero):

Hang on...

hero (hero):

a = original amount b = new amount \(\%\Delta = \) Percent Change \[\%\Delta = \frac{|a - b|}{a} \times 100\] \(= \dfrac{|11 - 33|}{11} \times 100\) \(=200\)

hero (hero):

So yes, you are correct.